OpenStudy (anonymous):

Using 4 equal-width intervals, show that the trapezoidal rule is the average of the upper and lower sum estimates for the integral (see below).

OpenStudy (anonymous):

\[\int\limits_{0}^{8} x ^{2}dx\]

OpenStudy (anonymous):

So we know the intervals are going to be 0,2,4,6 and 8

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

To find the trapezoidal area follow the formula: A trapezoidal=(Δx/2)(y0+2y1+2y2+2y3+2y4+y5) where Δx=(b-a)/n

OpenStudy (bohotness):

hold on love

OpenStudy (bohotness):

do you have answer choices?

OpenStudy (anonymous):

no

OpenStudy (michele_laino):

we have to compute the upper and lower sums

OpenStudy (michele_laino):

 that is the lower sum